### 19 Jan 2024 v1

# Rambler: Supporting Writing With Speech via LLM-Assisted Macro Revision

Abstract.
Dictation enables efficient text input on mobile devices. However, writing with speech can produce disfluent, wordy, and incoherent text and thus requires heavy post-processing. This paper presents Rambler, an LLM-powered graphical user interface that supports gist-level manipulation of dictated text with two main sets of functions: gist extraction and macro revision. Gist extraction generates keywords and summaries as anchors to support the review and interaction with spoken text. LLM-assisted macro revisions allow users to respeak, split, merge and transform dictated text without specifying precise editing locations. Together they pave the way for interactive dictation and revision that help close gaps between spontaneous spoken words and well-structured writing. In a comparative study with 12 participants performing verbal composition tasks, Rambler outperformed the baseline of a speech-to-text editor + ChatGPT, as it better facilitates iterative revisions with enhanced user control over the content while supporting surprisingly diverse user strategies.
copyright: nonejournalyear: 2024conference: CHI ’24: ACM CHI Conference on Human Factors in Computing Systems; May 11-16, 2024; Honolulu, Hawaii
Figure 1. Supporting long-form text composition via speech with a new dictation interface that consists of three components: 1) dictating trains of thought into Ramble boxes, 2) reviewing the spoken text through automatically generated keywords and summaries, and 3) performing manual or LLM-assisted macro-revisions operated at conceptual levels.
1. Introduction
Speech, regarded as a natural input modality, has attracted much attention following the breakthroughs of advanced speech recognition (ASR) and natural language processing (NLP). Speech transcription software products like Otter.ai, Google Voice Typing, and Dragon are gaining popularity. Dictation is now an integral keyboard function on most smartphones and tablets. Indeed, dictation can be three to four times faster than typing for text input in the best circumstances (Ruan et al., 2018). This makes it particularly advantageous for inputting long text, such as writing memos or drafting ideas. Mobile devices benefit even more from this given the challenges of using a keyboard and cursor on a touchscreen (Vogel and Baudisch, 2007).
Current interfaces for dictating long text remain largely the same as a traditional text editor with an added microphone. However, writing with speech can be more than just dictation. Previous research (Mehra et al., 2023) showed that speaking differs significantly from typing in terms of both production and memory retention. Transcribed spoken text can be much more verbose, erroneous, and disorganized, leading to high-effort editing (Karat et al., 1999). The verbatim memory of spoken content is also weaker, leading to a likely higher effort in reading the transcripts. Therefore, treating speaking simply as a faster “typing” method can be problematic, especially for long texts. Similarly, researchers for voice user interfaces (VUI) have made long-standing arguments against directly transplanting GUI guidelines to VUI design (Murad et al., 2019). All these call for new affordances and interface solutions for writing with speech. Meanwhile, writing is a highly iterative activity and thought development process. Typically, digital writing requires people’s full attention on a screen while sitting in front of a computer with a mouse and keyboard, but this requires time and an environment, which creates a high barrier to starting. While speech has the potential to be used for faster and lower-barrier writing, its adoption remains limited and the effort to edit resulting transcripts is large.
This work sets out to tackle that and answer the question: Can we identify and build stepping stones towards closing the gap between speaking and writing, by redesigning a dictation interface that caters to the shortcomings of speech? We build on existing solutions in recent works that use NLP and large language models (LLMs) to generate summaries to assist writing (Dang et al., 2022) and review spoken dialogues (Li et al., 2021, 2023a). These interfaces extract the gist of text to support visualization and/or interaction with it, which we call gist-based interfaces. They move beyond automatic text summarization to semantic manipulation, with the interaction centered more around the “gists” of the text rather than the sequences of characters of their words. While such concepts have been explored in traditional writing (Dang et al., 2022), we operationalize it in the context of interacting with spoken text, which is perhaps an even more compelling use case according to the literature showing spoken content remains in memory more as gists than verbatim (Mehra et al., 2023).
We introduce Rambler, an LLM-powered graphical user interface (GUI) that implements a gist-based interface tailored for dictated text (see Figure 1). Users can dictate their spontaneous thoughts and place each segment of recording (i.e., a fragment of an idea) into a “Ramble”. Rambler provides semantic zooming (Perlin and Fox, 1993) to visualize multiple summarization levels of gists for each Ramble, which users can review on demand. Users then focus on iterating on higher-level ideas by manipulating the Rambles, through respeaking, splitting, merging and reorganizing them, which we call macro revision. Behind the scenes, Rambler uses an LLM to automatically clean up the disfluency and punctuation errors from the raw transcript as well as to complete broken sentences and smoothen transitions, which we call micro revision. In addition, Rambler extracts keywords from the text to aid visual skimming across Rambles and semantic zoom levels. Individual keywords can also be activated and deactivated by the user to indicate importance, which then serve as parameters to customize the summary generation and other LLM-based word processing.
Rambler aims to aid the review and revision of spoken text while easing the interaction with text on mobile devices. To evaluate its effectiveness, we conducted a comparative study with 12 participants for a long-form verbal text composition task. The baseline is a standard speech-to-text editor interface with ChatGPT provided on the side. While participants demonstrated varied strategies and feature usage, most of them expressed preference of Rambler over the baseline and acceptance of using it for their own tasks. Their feedback shows Rambler provides better support for developing and reviewing spoken content, supporting iteration while improving user control over content revision.
Our contribution includes the following:
  1. (1)
    Rambler: a novel tool that implements the concept of a gist-based interface to support writing with speech on mobile devices. We motivate our design goals and provide implementation details for Rambler.
  2. (2)
    Empirical findings from a user study revealing the potential of Rambler in helping close the gap between rambling and writing. This study also identified advantages of our specialized LLM-backed GUI over a one-size-fit-all chatbot UI for LLMs, generating insights for designing and building LLM-supported direct manipulation interfaces.
2. Related Work
To provide more context for our contributions, we discuss three areas of highly related work: voice dictation and editing, supporting writers with LLMs, and semantic manipulations on text.
2.1. Voice Dictation and Editing
While speaking is faster than typing or writing, it can be time-consuming to correct errors caused by the speech artifacts, such as disfluencies and repetitions. Modern speech and NLP research has made much progress in cleaning disfluency and speech recognition errors (Liao et al., 2023; Lou and Johnson, 2020; Tanaka et al., 2018), mostly for post-processing. Researchers have proposed multimodal interaction methods to reduce editing effort, especially on mobile devices. For example, EyeSayCorrect uses gaze to infer users’ editing intentions and allows re-dictation to edit the phrase near eye fixations (Zhao et al., 2022). Similar ideas were applied to other modalities, such as touch (Zhao et al., 2021).
Other work has explored speech-based editing through automatic removal of colloquial or filler words (Arnold et al., 2021) as well as voice editing with and without explicit command keywords (Ghosh et al., 2020). Ghosh et al. (Ghosh et al., 2020) found that using explicit command words (e.g., ”replace” and ”delete”) were suitable for one-word edits, and re-dictation without command words was suitable for complex edits. To leverage the advantages of both cases, ”Just speak it” supports commanding without explicit command words (Fan et al., 2021), which are instead induced from existing context and the edit command. Li et al. (Li et al., 2023b) further explored auto segmentation and classification to support the seamless transition between voice dictation and editing with natural language. While this line of work lays the groundwork for natural and effortless voice-based text editing, it primarily focuses on editing one or a few words. Generalizing to sentence-level macro revision remains largely unexplored.
2.2. Writing with LLMs
A number of studies have demonstrated that text generation by LLMs can be beneficial for writing assistance. For example, Yuan et al. (Yuan et al., 2022) developed Wordcraft, an editor that allows users to collaborate with an LLM to write stories. Users can replace selected text with LLM suggestions or prompt the model for text generation. Yang et al. (Yang et al., 2022) investigated an LLM-powered writing tool to shorten, edit, summarize, and generate text. Gero et al. (Gero et al., 2022) customized the decoding method of an LLM to encourage diverse outputs specifically for scientific writings.
While LLMs have made significant progress in the field of writing, there is still room for improvement. LLMs may struggle to retain the writing style and align with the writer’s goal. Ippolito et al. (Ippolito et al., 2022) found that LLMs can struggle to preserve the style and authorial voice of experienced writers. This is supported by Biermann et al.’s (Biermann et al., 2022) study, which found that authors prefer text generation tools that respect their styles and strategies. Rahman et al. (Rahman et al., 2023) studied academic writing with ChatGPT. Their findings revealed that ChatGPT can be useful to generate initial ideas, but falls short in tasks that demand critical thinking, such as literature synthesis, citations, problem statements, research gaps, and data analysis. Taken together, LLMs are a promising tool for writing, but they alone are not yet perfect.
2.3. Semantic Manipulations on Text
In addition to text generation discussed in the previous section, LLMs can also support macro-level structural revision. However, as Strobl et al. (Strobl et al., 2019) recently surveyed, writing tools that support grammar and word-level editing are common, but tools for macro-level structural revision are relatively rare. Arnold et al. (Arnold et al., 2021) highlighted the opportunities to return semantic control to the writers. One of their proposals is to support within-sentence structural manipulation, where users can drag and drop words to rephrase the sentence. Alongside that, semantic zooming (Perlin and Fox, 1993) is another promising feature for collaborative human-AI interfaces. It allows users to explore and navigate complicated data (e.g. long texts and nested tables) by zooming in and out to check details of different levels. This can be particularly useful for interacting with LLMs in a conversational paradigm, as it reduces the cognitive and physical load of scrolling back and forth to linearly search for information from a lengthy conversation. Several studies have demonstrated that semantic zooming is an effective way to help users navigate and digest complex content. Sensescape (Suh et al., 2023) and Graphologue (Jiang et al., 2023) introduce zoomable visual representations of different concepts involved in a conversation with LLMs. Li et al. (Li et al., 2021, 2023a) developed a system that generates hierarchical summaries of spoken dialog. As the closest work to ours, Beyond Text Generation (Dang et al., 2022) is a writing tool that provides on-the-fly paragraph summarization along with the original text. This work also provides interaction with the summaries of paragraphs, such as reorganization via drag and drop, which manipulates the original text in parallel. Building on this work, we leverage the benefit of summaries as one way of extracting gists. While its focus is on keyboard typing and ours is on speech input, our interface also provides keywords extraction and interaction as well as other LLM-assisted operations in order to support macro revision.
3. Rambler: Gist-based Interface for Writing With Speech
3.1. Design Rationale and Goals
As explained in the introduction, we aim to design a gist-based dictation interface to support writing with speech by helping users overcome the shortcomings of speech production and memory. The shortcomings identified from previous empirical and theoretical works (Liu et al., 2022; Mehra et al., 2023) can be summarized as follows: 1) speech is spontaneous as it is produced in real-time, so while the gists may be planned, the wording is not; 2) looking at the live transcript while speaking can be distracting, so people tend to have less visual engagement speaking than they would typing, which leads to less micro-revision when speaking; 3) the verbatim memory of spoken content is weaker, which, combined with less visual engagement, may make reviewing content harder.
Figure 2. A labeled screenshot of the Rambler UI. (1) Ramble in default state, with macro-revision interactions accessible through buttons on (2) the top left and (3) the top right. (4) Ramble in re-speaking mode, where voice input is transcribed so that it can be appended to current text, replace the current text, or be discarded using the buttons from left to right at (5). Other macro-revision features are temporarily inaccessible. Fixed at the bottom of the UI is (6), with the semantic merge button, new Ramble button, and semantic zoom slider.
3.1.1. Rambles as the unit of interaction
With these characteristics, a text editor that transcribes everything linearly will also include all the disfluencies, repetition, and disorganization of raw speech. Although advanced NLP techniques can help correct some of these issues, the user still needs to direct the composition’s meaning and organization. Successful LLM interfaces (e.g.,ChatGPT [^1] https://chat.openai.com/, Claude [^2] https://claude.ai/) have demonstrated the ability to refine unorganized text into high-quality writing. However, such general tools not only rely on users’ ability to prompt the LLM, which was observed to be challenging (Zamfirescu-Pereira et al., 2023b, a), but are also generally hard for users to precisely control where and how their text is edited.
Building on previous works that show the benefits of using NLP or LLM in direct manipulation (Dang et al., 2022; Jiang et al., 2023), we investigate how to leverage recent LLM and NLP techniques to inject intelligent word processing into the GUI for our specific task of writing with speech, with the goal of preserving user control over the their composition. With these considerations, we aim to improve the text editor interface for better interaction with spoken text by establishing the following design goals:
  • Support non-linear composition. The tool should allow users to ramble spontaneously while providing them with structures to support iterative organizing and editing their spoken text.
  • Support iterative drafting. Leverage the fast production of speech while recognizing that the first attempt may not be well articulated as written text. Our interface should encourage and support users to easily iterate on drafts by respeaking large chunks of the text before diving into micro-editing.
  • Aid the review and navigation of spoken text. The interface should help reduce user effort in reading and comprehending the transcript, which could be verbose, repetitive, and contain errors or disfluencies.
  • Leverage LLM while preserving user control. The interface should leverage the capacity of LLM in cleaning up, smoothing, and transforming text while allowing users to control where and how the change will be applied.
  • Optimize interactions for mobile devices. Although speech-for-text input can be used on any device, mobile devices particularly benefit from it, given the known challenges of interacting with text on these devices without keyboards. For our evaluation, we target a tablet interface for designing our current interface, as one of the pervasively used mobile devices suitable for long-form writing.
In the next subsection, we will explain how we designed the interface of our tool to achieve these design goals.
3.2. Rambler Interface Design
Here, we explain the design of Rambler and its main features supporting macro revision and text reviewing, respectively.
Rambler is built around the creation and manipulation of “Rambles”, illustrated in Figure 1 and 2. Each Ramble is a container (visually separated from other Rambles) that holds its content and multiple buttons for users to interact with. As we envision it, the content of a Ramble roughly corresponds to a single train of thought. Without the pressure of producing a decent piece of text in one go, users may either speak everything into one Ramble and split it at any time, or start a new Ramble when they feel what they will say next is about a different point or subject matter. This should support various writing strategies with or without an outline in mind.
3.2.1. Supporting text reviewing with clean transcript and gist extraction
Rambler is designed to help users review the dictated text more efficiently with three key features (transcript cleaning, semantic zooming, and keyword spotting) that leverage NLP techniques and LLMs.
Figure 3. From left to right, example text in three Rambles is presented at all four semantic zoom levels: full transcript, 50% length, 25% length, and 10% length.
Figure 4. Example Rambles before and after (A) semantic merging and (B) semantic splitting. In (A), note that the user must tap to select all the Rambles they wish to include, then execute the semantic merge with the global button (visualized in Figure 2). In (B), the user taps the scissors icon in a Ramble to automatically split it.
Figure 5. Example Ramble showing (a) keyword spotting, with keywords highlighted in light green, and (b) the magic custom prompt modal, with an example prompt.
First, LLMs are used to automatically clean Ramble transcripts when they are added or edited, in order to minimize small speech-to-text (STT) typos (like messy punctuation or misheard words) that impact reading. To avoid distraction, our interface only shows raw speech for real-time feedback when the user is actively dictating; if the user stops dictating, the interface immediately cleans the spoken text, and only displays the clean transcript.
Second, we implement semantic zooming through a slider that controls the degree of summarization at which we view all the Rambles in the composition. As shown in Figure 2.6 and Figure 3, users can move the slider at the bottom of the UI to view text summaries at different levels, specifically at 50%, 25% and 10% of the original length. [^3] Rambler’s prompts for these summary levels were tuned to achieve these levels, but are not explicit to these percentages; see A.2.4. We design this feature to help visualize the core ideas of lengthy text and to allow users to quickly gain a sense of how their ideas should be re-organized and iterated upon.
Last, to aid visual skimming across Rambles, Rambler automatically highlights keywords after a Ramble is added or edited (see Figure 4(a)). Words in the Ramble can be toggled by the user with a tap to be selected/deselected as a keyword. Besides being a visual aid for continuity across zoom levels, users can also tap a regenerate button on a Ramble to re-summarize using their transcript text and their modified keyword selection. This is to address an issue we discovered that summaries generated by LLMs may not be aligned with the users’ idea of Ramble content.
3.2.2. Supporting macro revision with LLM assistance
Figure 6. Rambler System Architecture Diagram. (A) The user provides speech input to be processed or provides touch input to the interface. (B) Rambler user interface. (C) Our cloud server mediates between user requests and the OpenAI API. (D) The AssemblyAI API is used for transcription and OpenAI GPT-4 API for LLM functionality.
We distinguish between conceptual-level edits, which we call macro-revisions, from fine-grained edits, which we call micro-revisions. Macro-revisions operate on ideas, and in our prototypes happen at a Ramble level – consider functionalities like reorganizing Rambles, and merging or splitting Rambles. Micro-revisions, in contrast, operate on a much smaller scale and, in our prototypes, happen primarily through keyboard editing – consider fixing a single typo or some word choice in a sentence. We consider macro-revisions and micro-revisions complementary, thus our interface also provides keyboard editing, activated by clicking “Edit Text.”
We designed five macro-revision operations: Ramble respeaking, Ramble re-ordering through drag and drop, Ramble merging (concatenation of text) by dragging one Ramble onto another, Ramble splitting (by pressing return during keyboard editing) and Ramble transformation. Although macro-revisions can also be done in a normal text editor, we provide an interface with stronger affordance for doing so, with the goal of encouraging and supporting users to apply these operations on spoken text. We carefully designed a re-speaking mechanism and interface (Figure 2.4), which is activated when the user re-activates the mic of an existing Ramble. The user talks further and then chooses whether to append the new text to the Ramble, replace the old text in the Ramble with it, or discard the new text (Figure 2.4). During the respeaking process, the original text remains above in grey color as a reminder. We do so to aid the user’s memory, as in our design process we learned that without the original text, it is hard to remember what to say.
While all of the macro revisions can be performed manually, we provide an LLM-assisted version for three of the operations – Ramble merge, split, and transformation – and named them semantic merge, semantic split and magic custom prompt.
Semantic merging utilizes the LLM to intelligently merge the content of multiple selected Rambles into the content of one Ramble by clicking a “Semantic Merge” Button (see Figure 2.6). The semantic split button is placed on individual Rambles, which prompts the LLM to divide one particular Ramble into N Rambles based on its content. These two mechanisms are visualized in Figure 4 with example text. The magic custom prompt function is triggered by a magic wand button on a Ramble. It opens up the possibility for expert LLM users to define any custom transformation on Ramble content by directly inputting an LLM prompt (see Figure 4(b)).
3.3. Example User Workflow
To better illustrate the functionality of Rambler, we consider an example of a simple user workflow: Alex, an artist, is drafting a post for his personal blog and wants to describe how he finds creative inspiration. He has many ideas on the topic, but isn’t sure the best way to write it: hence, he reasons that he should simply talk it out. Opening up Rambler, Alex presses the + Ramble button to begin recording. After finishing his thoughts and stopping the recording, Rambler automatically cleans this into the first Ramble. Alex senses that he wandered somewhat from his initial topic; since the resulting text is fairly long, he slides the semantic zoom slider to see a summarized version of his text. Then he takes a look at the highlighted keywords in all the zoom levels to get a better sense of what he talked about. After glancing back at the transcript, Alex decides to utilize semantic splitting to revise the various ideas in his Ramble separately.
Rambler segments his cleaned spoken text into multiple rambles along three main ideas, 1) ”finding personal inspiration”, 2) ”interaction between self expression and audience interpretation”, and 3) ”difference between art and design.” With a clearer understanding of the ideas he’s just verbalized, Alex begins to alternate between composing new related Rambles, and revising with the provided functionalities (eg. respeaking, reordering, merging, and/or splitting Rambles) – until he feels like each Ramble roughly resembles a somewhat concrete idea. Once he feels that he has added all his thoughts and developed them into a cohesive flow of ideas, he uses the built-in keyboard to fix word choices here and there to his liking, and then exports the composition to share with the world.
3.4. Prototype Implementation
Rambler is implemented as a React.js web application with a node.js backend and is hosted in the cloud. LLM-based text manipulation functionality is enabled using OpenAI’s GPT-4 API, which provides high quality output for our desired LLM functionality. We considered other models such as GPT-3.5 Turbo for response time reasons; their output on LLM interactions (most notably, semantic split) was less precise. Rambler’s voice transcription uses AssemblyAI’s real time Speech-to-Text API. We made this engineering choice because AssemblyAI’s API allowed us to transcribe text indefinitely in a single instance. Automatic keyword selection is implemented via the rapid automatic keyword extraction (RAKE) method (Rose et al., 2010). Compared to an LLM based approach for keyword selection akin the method used in Graphologue (Jiang et al., 2023), the RAKE method was orders of magnitude faster ($\sim$4s vs nearly instant). Our overall system architecture is illustrated with Figure 6. Our project source code is available at [LINK ANONYMIZED] and a demo can be found at [LINK ANONYMIZED].
GPT-4 was used for the four key LLM-powered gist-related features: text summarization for semantic zoom (including transcription cleaning), merging multiple Rambles, splitting a single Ramble into multiple, and open-ended editing using customized LLM prompts. We created pre-defined prompts for zoom, merging, and splitting; the open-ended editing prompts are entirely user defined. One-shot prompting was employed for the transcription cleaning prompt and zero-shot was used for the rest. Keywords are appended to prompts when marked for inclusion through selection. All prompts are attached in Appendix A.2 for reference.
4. Evaluation
To evaluate the effectiveness of Rambler for supporting writing with speech, we designed a within-subjects study to compare Rambler with a state-of-the-art solution for dictation and word processing. The baseline is chosen to be: raw dictation into a text editor with ChatGPT provided on the side. Participants completed two long-form writing tasks through dictation, one using Rambler’s full suite of available functionality and the other using the baseline. Through this comparison, we seek answers to the following questions:
  1. (1)
    Can writing with speech be better supported by Rambler than the baseline? If so, how?
  2. (2)
    How effectively does our approach of embedding LLMs in GUI support the review and revision of spoken text?
4.1. Baseline Interface
As mentioned above, the baseline uses a plain text editor interface plus dictation. Raw transcripts go into the text box as a user speaks after the microphone is activated. The transcript can be edited using iOS’s built-in keyboard and cursor interface. To avoid noises caused by different speech recognition engines, we implemented this interface using the same speech recognition engine (Assembly.ai) used in our prototype as a web interface. To make a fair comparison between the baseline and Rambler, ChatGPT is provided with a separate tab in the same web browser. However, we limited the use of ChatGPT to revising content produced by the participants by not allowing them to ask ChatGPT to produce major compositions for them.
4.2. Participants
Twelve participants ($N=12$) were recruited from our university campus community through various mailing lists and student organizations. All participants were native-fluent English speakers and there was varied experience with dictation and LLM use. P3 and P10 were aged between 25 and 34; all other participants were aged between 18 and 24. No participant had motor impairments that impacted keyboard or dictation usage. Further participant details are presented in Table 1.
4.3. Experiment Task
The task for the experiment is to compose and revise a blog post using the given interface in each condition. The reason for this task setting is to encourage participants to compose long-form personal writing and revise it extensively for an audience, in order to combat the potential lack of motivation to improve the text in a lab study. Participants were asked to choose a topic they were personally interested in, from a provided list of open-ended writing tasks (e.g., “Think of two dream jobs of yours.”, “Think of two academic concepts you find interesting.”, and so on). Then, for each of the two conditions, they were asked to create a “blog post that would be posted on your social media” about one chosen answer to the topic. Notice that we had participants pick two variants for the same topic (such as two dream jobs instead of a dream job and an academic concept) to minimize the influence that completely different topics would have on participants’ compositions.
4.4. Procedure
A lab study session for each participant lasted about 90 minutes, which could be divided into three main parts – the pre-study survey, the main study, and the post-study interview. The study session was conducted in a quiet room in person with an experimenter. Participants accessed Rambler via Safari on an 11-inch iPad Pro running iOS 16. There was no hardware keyboard. Nonetheless, the participants were allowed to use the on-screen keyboard when needed.
The pre-study survey.
Participants were first given a brief, initial demographic survey including questions about their familiarity with writing, dictation tools, voice-interface tools, and LLMs.
The main study.
Participants were not told we made the interfaces. In a counterbalanced order across participants, each participant used both two conditions. For each condition, the participants engaged in 1) a training trial, 2) the task, and 3) a post-task survey. In the training trial, the interviewer spent about 5 to 15 minutes explaining all the functionality of the given condition and provided time for the participant to try and familiarize themselves with each feature. Next, to perform the task, the participants were given up to 20 minutes to improve their composition until the time is up, or until they felt satisfied, whichever came first. If the participants hit a word count of 800 words or more, or had not done any editing by the 10-minute mark, the experimenter would encourage the participant to begin editing (so that they have a decent amount of time to revise their composition). In the post-task survey, the participants were asked to assess the usefulness of each feature in the given condition using a five-point Likert scale. They were also asked to estimate a percentage on how much micro-revision (specifically keyboard-editing) they did in comparison to macro-revisions. They were then asked to assess the quality of their text composition on a five-point Likert scale.
Table 1. Demography and dictation/LLM usage of study participants.
ID Degree Level Domain Dictation Usage LLM Usage
P1
Ph.D.
Computer Science
A few times a year
A few times a week
P2
Undergrad
Computer Science
Never
Everyday
P3
Ph.D.
Computer Science
Never
Never
P4
Undergrad
Computer Science
A few times a month
A few times a week
P5
Undergrad
Business
A few times a week
A few times a month
P6
Undergrad
Cognitive Science
Never
A few times a week
P7
Ph.D.
Electrical Engineering
Never
A few times a year
P8
Undergrad
Data Science
Never
A few times a month
P9
Undergrad
English
A few times a month
A few times a month
P10
Ph.D.
Applied Physics
A few times a year
A few times a month
P11
Undergrad
English
A few times a year
Never
P12
Undergrad
Computer Science
Never
Never
The post-study interview.
Following the two conditions, the participants were interviewed for about 25 minutes regarding their overall experience. Specifically, they were asked about their workflows for each condition and their experience with the features in both conditions. Finally, the participants were asked for their overall preference between the baseline and the Rambler condition. Each participant was compensated \$30 USD for the study.
4.5. Data Collection and Analysis
We captured three types of data for analysis: a screen recording of the iPad running the Rambler application; audio recording of participant and interviewer speech during each study session; and application logs of user interactions with the interfaces. We performed descriptive quantitative analysis of the ratings and logs, and summarized our interview findings in qualitative analysis.
Qualitatively, we engaged in exploratory data analysis, transcribing all audio recordings, and coding the screen-recording videos to observe participants’ workflows and strategies. Two of the authors then compared these approaches across participants and categorized their usage through a thematic analysis, relying on a modified form of affinity diagrams (Moggridge, 2006) and service blueprints (Bitner et al., 2008) to document the specific workflows and strategies each participant engaged in and the experiences they communicated with the interviewer.
5. Results
This section synthesizes the findings from our quantitative and qualitative analysis. We analyzed the overall workflow and identified a broad diversity of ways Rambler was used. The questionnaires and interviews compared the effectiveness of Rambler in supporting writing with speech, and collected feedback on user acceptance. The usage of each feature in Rambler was reported in reference to logs, ratings and interviews.
5.1. Overall Workflow, Strategy, and Outcome
While this section summarizes and highlights interesting findings about participants’ workflows and strategies, a detailed visualization figure of operations performed in all the tasks is shown in Appendix A.1.
5.1.1. Workflow
Participants adopted different workflow strategies with the baseline method and Rambler, with Rambler encouraging more iterative writing throughout the writing process. In the baseline condition, seven of 12 participants dictated most of their text first before doing more in-depth revision, while five alternated between composing through dictation and revision. Nine participants explained that they put the dictated text into ChatGPT before doing any keyboard editing, as the baseline transcript had extraneous punctuation and occasional word recognition errors. In the Rambler condition, four of 12 participants composed the bulk of their text upfront and then spent the rest of their time in revision, while eight moved back and forth between composition and revision multiple times in the task.
5.1.2. Composition and Revision Strategies
In the baseline condition, nine participants first ran their text through ChatGPT. They mainly asked ChatGPT to: clean up the text, improve clarity, change the tone or style (e.g. to be like a blog post, to sound like an introduction or conclusion, or to sound more formal). They came up with prompts like: ”Clean up this transcript, fixing any typos”, ”Can you remove any repetition from the text?”, and ”Format this into a more casual blog style article.” Three participants did not take this route, instead starting with keyboard editing. Interviews revealed that the main reason was a lack of trust in ChatGPT to not add unwanted content or to generate something aligned with the personal story.
With Rambler, participants exhibited a large diversity of approaches in using the various features for splitting, merging, and keyword use offered in Rambler; no single consistent workflow emerged as the most common. Five participants (P1, P5, P7, P9, and P11) developed revision patterns (e.g., keyboard edit and magic custom prompt) that they would use for all the rambles before doing a composition-level review. Six participants (P2, P3, P4, P6, P10, and P12) pursued more restructuring through select combinations of macro-revisions before keyboard editing. P8 had a more unique approach: they indirectly edited the 50% semantic zoom summary to generate their final composition. The list below highlights the distinctive and creative high-level strategies used by each participant. Refer to the timeline visualization in Figure 10 in Appendix A.1 for better comprehension.
  • P1 typically started new rambles with a few words, re-speaking to add on, and then keyboard editing it. At the very end, they manually merged all rambles together to apply magic custom prompt, and then they used semantic split to get paragraph breaks.
  • P2 thought about what to say before actually starting speaking out loud, by pre-dividing their thoughts into Rambles. They then edited the composition at a higher level through semantic merges and the magic custom prompt, before diving into more keyboard editing.
  • P3 used Rambles like subpoints, creating an outline for themselves. They recorded a couple of rambles, used a series of magic custom prompts to clean the entire composition, recorded more rambles, then used semantic merge and manual split to restructure the composition
  • P4 first produced a short paragraph and used magic custom prompt to reformat it into a design document (which inherently generated more text) and then used semantic split and manual merge as they continued to add content through speech. They included more keyboard editing in the latter half of the task.
  • P5 did not pre-plan, and spoke as if telling a story; they keyboard-edited each section without much macro revision, but at the end merged everything into a single Ramble to apply the magic custom prompt “clean up text”.
  • P6 added content as it came to mind, and reorganized, manually merged and split, applied magic custom prompt to revise tone, and keyboard edited text.
  • P7 spoke in a sequential way—for each ramble, they would speak, then keyboard edit, and occasionally use magic custom prompt to help revise. At the very end they took a look at the semantic zoom summaries.
  • P8 recorded most of their rambles first, then visited 50% semantic zoom level to see what was there. They indirectly edited the 50% zoom level by regenerating summaries, semantically merging, and adding onto rambles until they were happy with the text in the 50% zoom level summary, which they used for submission.
  • P9 worked in a fairly linear order. After recording each ramble, they would sometimes choose to keyboard edit or alter keywords, and then use magic custom prompt (with keywords as context) to “clean the text”. At the very end, they did some reordering.
  • P10 recorded all their content first, then highlighted keywords in all the rambles and regenerated all the summaries. Then, they semantically merged all of them together to see what this would produce; because the context shifted due to LLM variability, they had to do some keyboard editing at the end.
  • P11 recorded all of their content first, and then keyboard edited before using magic custom prompt (using keywords as context). At the very end, they took a look at the semantic zoom summaries.
  • P12 rambled a lot to branch off ideas and then restructured using reordering, semantic split and merge, and manual merge. They highlighted keywords for their own reference.
Figure 7. Final text quality Likert rating by condition (technique)—Baseline and Rambler. The left side shows the individual ratings and the right side is a boxplot summarizing the comparison.
5.1.3. Text Output Quality
We asked participants to rate the quality of the text they submitted as final outcome for each task. As shown in Figure 7, more than half of the participants rated the outcome quality the same across both conditions. Two participants rated higher for the baseline and three rated higher for Rambler. An independent t-test of the Likert ratings revealed no significant difference in perceived text quality across conditions, consistent with the overview we get from the boxplot in Figure 7. (See Section8 for a discussion of assessing text output quality.)
Figure 8. Stacked bar graph of feature usage counts associated with (a) text reviewing and (b) macro-revision. In (a), one usage count for semantic zoom represents a single transition from a semantic zoom level to an adjacent level (e.g. a drag of the slider from 50% length summary to 25% length summary). All other usage counts in (a) and (b) represent a single tap, drag, etc. on their respective interfaces.
5.2. Effectiveness of Rambler Compared to Baseline
The questionnaire and interview reveal that Rambler is an effective tool for supporting text composition and revision via speech, especially for tasks that require organizing one’s thoughts, iteration, and maintaining one’s writing voice. We explicitly asked participants about Rambler with respect to reviewing composition, organizing thoughts, supporting iteration, and granting user control.
5.2.1. Support for Reviewing the Composition
Eight out of 12 participants found Rambler to be more helpful for reviewing text composition than the baseline system. Half of these participants attributed this to the Ramble structure, which made it easier to discretize, visualize, and review their text composition. The other half found semantic zoom to be useful for grasping the main idea of their Rambles and checking that they included or emphasized all the points they wanted to. Additionally, the automatically generated clean transcript made it easier for participants to review their text.
5.2.2. Support for Organizing Thoughts
Eleven out of 12 participants felt that Rambler was better for organizing their thought process. Most participants attributed this to the ramble structure and its inherent support for reorganization. For example, P6 said that the Ramble structure made it ”a lot easier… to keep track of what I was doing.” P1 explicitly said that Rambler was ”clearly better for organizing… I was able to move the ideas around. And that was very helpful.” P4 also found it ”super useful for dragging my thoughts around and kind of restructuring.” P3 was the only participant who felt that the baseline was better for organizing their thought process. They rationalized this by the free-form nature of the baseline’s text editor. Their preferred workflow depends on a nested structure which Rambler doesn’t support.
5.2.3. Support for Iteration
Seven out of 12 participants felt that they iterated more in Rambler. This was due to the Ramble structure, the support for macro-revisions, and the automatically generated clean transcript. The Ramble structure made it easier to discretize text composition and allowed for more flexibility, which made iteration faster. P3 said that Rambler was ”a lot more …inviting of … merging and smashing things together. So I could definitely iterate a lot more.” The clean transcript made it easier to review on the spot and enabled iteration on the go. P10 and P11 explicitly mentioned that they iterated more in Rambler because they did not feel motivated to iterate at all in the baseline.
5.2.4. User Control
10 out of 12 participants felt that Rambler gave users more control over the content iteration. While some functionalities like manual merging, manual splitting, and reorganization as pointed out could be mimicked in the baseline, five participants emphasized how the support within Rambler made it faster to perform macro-revisions. Knowing that there was support for iteration made participants feel more in control. Furthermore, the structure of the Rambles helped localize changes. Four participants mentioned that this made it easier to know what part of the text the intelligent text processing was going to be operating on. P8 said that ”when you figure out what [Rambler ] does, it is much easier to really customize your composition… If you want more detail, Rambler is worth learning.” Some users also noted how they felt more in control of their writing voice. P11 and P12 were pleased at how the clean transcript more accurately preserved their voice than ChatGPT could. P11 explicitly mentioned how keywords helped preserve their voice when they utilized the LLM through Magic Custom Prompt. Among many other participants, they expressed that Rambler made it possible for them to “see it better and edit it better”.
5.3. Usage of Rambler Functions
Here, we analyze participants’ feature usages from both quantitative and qualitative data. Overall, the participants in the study found Rambler to be a useful tool for writing. As detailed in Section 5.1.2, participants invented diverse and creative strategies when using Rambler. These diverse strategies led to equally diverse usage of functionalities. Figure 8 visualizes the count of usage of each function logged in our system, per participant. As we can see from Figure 7(a), semantic zoom and keyword spotting were the most frequently used features. Overall, the use of features varied largely across participants. Figure 9 shows the subjective ratings of each function in the post-task survey. The following analysis assesses the use of each feature by referring to these ratings and summarizing relevant insights revealed from the interview. The most well-received feature is the Rambles structure. Some LLM-assisted operations were less commonly used, but participants still found them to be very helpful.
5.3.1. The Ramble Structure
Participants made good use of the Rambles structure to organize and discretize their writing. While they had varying opinions on how it impacted their writing process, all 12 participants found it useful for its organizational benefits. Some participants found the structure more efficient when they had ideas prepared beforehand, while others preferred it for freely brainstorming. Overall, participants found that Rambles helped them to better discretize their composition, which made it easier to visualize and revise. Specifically, eight participants mentioned that Rambles helped them to focus on and iterate on specific sections of their work. P3 and P5 appreciated that the sections of Rambles mimicked their usual writing style of outlining first, which helped them to get used to the structure. P12 noted that the structure allowed them to separate their ideas and feel more in control of their organization, while P9 found that it enabled them to work in a more methodical manner.
Figure 9. Usefulness Likert Rating per Feature in Rambler. For each feature we had the statement, “I found $<$feature_name$>$ useful,“ and asked participants to indicate whether they “Strongly Disagreed“, “Disagreed“, felt “Neutral“, “Agreed“, or “Strongly Agreed.“ The total number of ratings per feature varies because some participants indicated that they did not use a particular feature.
5.3.2. Manual Operation Usage for Macro Revision
Rambler offers a variety of manual operations to help users revise their writing in larger pieces, including respeaking, reorganizing, manual merge, and manual split. Participants generally found them helpful, while individual feature usage differ based on their workflow.
Respeaking
Four participants used respeaking to replace entire Rambles. A small number of participants also used respeaking to add new text to a Ramble, sometimes splitting a particular sentence out of a Ramble into a new one to replace that sentence with an expanded dictation. Half of the participants found this feature to be a nice addition, and more than half of the participants agreed or strongly agreed with the statement that respeaking was useful in their ratings. While some mentioned that they could simulate it by deleting and restarting the Ramble box, one participant found that respeaking is quicker and more convenient.
Reorganizing
Participants’ usage of reorganizing Rambles depended on their individual workflows. Some participants (three out of 12) did not use reorganizing at all, while four others said that they did not reorder their compositions too often. However, most participants felt that reorganizing was a useful feature, which is supported in the Likert ratings, even if they did not use it frequently. Interestingly, a couple participants felt more comfortable rambling knowing that they could easily adjust their ideas afterwards. As P4 put it, “I was able to just kind of get my ideas on the screen without necessarily like a linear progression, because I knew I could… move things around [after]… the Rambles… [help let] me… freely brainstorm, but then allowing me to kind of put it all together in a way that made sense.” While several participants pointed out this can also be done in a normal editor by copy and paste, P12 noticed in a “mobile device [context]…it was a lot easier… copying and pasting text [in the baseline]… was a little bit more clunky”.
Manual Merge and Split
Manual merge and manual split were used by participants depending on their individual workflows. While many participants found manual splitting and merging to be helpful and useful as further indicated in the Likert ratings, these two features were not considered essential by all. Some participants used them to combine Rambles, take breaks in between dictation, or section out parts of the text for respeaking. P3 utilized manual splitting as a way for them to regain control over and “rein in” their composition after using semantic merge. Some participants also found tension between manual splitting and merging and semantic splitting and merging, as they preferred one or the other depending on their workflow.
5.3.3. Semantic Operation Usage for Macro Revision
Rambler provides a number of LLM-assisted operations, some of which are more well-received than others. Cleaning transcript is a valuable feature that the majority of participants found helpful. Semantic zoom is a versatile feature that can be used for a variety of purposes, but it is not as widely used as part of people’s strategies. Semantic merge and manual split are less commonly used, and some participants stopped using them after a few attempts.
Cleaning Transcripts
Most participants found the live cleaning of the raw transcript to be helpful and useful. P4 and P12 specifically mentioned how much they appreciated the feature after using the baseline (where they didn’t have the automatic clean transcript). Participants P11 and P12 were also pleased at how the clean transcript preserved their voice better than what they could get using ChatGPT in the baseline. The accuracy of the clean transcript also made it easier for participants to immediately and actively review, understand, and iterate on their composition. As P11 expressed, the cleaned transcript was so high quality that they were willing to do keyboard edits to it (whereas in baseline they just handed it off to ChatGPT).
Semantic Zoom
The usage of semantic zoom varied depending on each participant’s strategy. Most participants used or explored using this feature according to Fig. 7(a). It was used for a variety of purposes, such as gauging whether they could query the LLM for more succinctness (P4), review what they talked about (P7, P8, P11), checking if they missed any main ideas (P7), using generated summaries as part of final text (P8), or simply out of curiosity P10. The relatively high count of usage in Fig. 7(a) is partially due to the way logs were generated - if a participant scrolls through the slider it gets registered at each level change. Semantic zoom was observed to be less used when participants did more keyboard editing, perhaps because participants can gain a good understanding of the text while micro-editing and thus need less assistance for reviewing. P3 found semantic zoom less helpful because their composition was short enough to not need summaries and felt that the shortest level was too summarized to be useful. On the other hand, P7 was very impressed with it, and P8, who used semantic zoom as a tool for editing instead of reviewing, expressed that it was the most useful tool of them all. P8 used the 50% level as the final submission because it felt like a “more clean or more postable version” of their work, and mentioned how they “liked having more information to go off when I’m reviewing my work.” They also appreciated how easy it was to regenerate summaries by customizing keywords.
Semantic (LLM-assisted) Merge and Split
The usage of semantic merge and split varied depending on participants’ workflows. Seven out of 12 participants used semantic merge and three used semantic split. Participants used semantic split to partition long Rambles (P4), and participants used semantic merge to achieve cohesive combinations (P2 and P4) or out of curiosity (P3, P10). P3 mentioned it is good for low-stakes where word choice precision doesn’t matter as much. When it comes to the reasons why semantic split was less used, we observed that participants used manual split more often than semantic split, perhaps because it warrants more certainty. As P3 said, “I sort of use the manual splitting as a way to like rein in the semantic merge, which I really liked.” Some (P7, P9) felt the task wasn’t long enough for semantic split to be useful. P10 was worried about unexpected results thus was reluctant to use. We also learned that some participants stopped using these features after a few attempts because the features gave unexpected results (P1).
Magic Custom Prompt
Most participants made extensive use of the magic wand tool to revise within Rambles: change the tone of, reformat the structure of, further clean up their text composition, and even reducing the size of a Ramble. One participant spoke with the intention of constructing an outline, then used the magic wand tool to expand this outline into a “design document” (P4), which they then split into multiple Rambles for further revision. Four participants had positive outlooks on the magic custom prompt, though P2 felt that they were relying on it too much. P11 and P12 also liked magic custom prompt, but primarily because it preserved their writing voice more than ChatGPT in the baseline did when they used keywords to influence the outcome. In contrast, P8 did not like this feature, as they felt that it was not helpful in getting the style or language that they wanted.
Keyword Spotting and Regenerating Summaries
Keyword spotting (and regenerating summaries) was used in different ways by participants, but those who adopted it agreed or strongly agreed it was useful. Participants used keywords in a variety of ways: to understand their readers (P7), improve their focus and understanding of their own writing (P2, P12), highlight keywords to regenerate summaries (P8, P10), and influence their magic custom prompt (P9, P11). P4 found that highlighting keywords could be extra work, as they could change during editing. P11, on the other hand, found that providing keywords as context to the system helped them retain their writing voice.
5.4. User Acceptance and Use Cases
5.4.1. On using speech to write
Dictation is a powerful tool for capturing thoughts for writing, but it can take time to learn how to use it effectively. Our post-study interviews revealed mixed opinions on the experience of writing with speech, indicating a learning curve to overcome. Three participants who used it a few times a month or year (P9, P10, and P11) found it to be “a really good way” to get started and “capture [your] loose thoughts before they run away” (P9). P4 (a few times a month) felt both tensions: it “kind of just let your ideas flow” to use dictation, but there’s also “a little build up in my brain of… what should I say first.” One participant who reportedly used dictation weekly (P5) described their process as “I speak before I think” but “that’s not exactly a good thing” because their “words are definitely not clean”— and that Rambler encouraged them to “ramble a lot more”. Finally, participants with no experience (three who had never used it before: P2, P3, and P7) felt that speaking made them think about what to say first and dictation slowed them down. (P6) also mentioned that they did not like having to manually clean up STT transcripts. Overall we did not find a consistent trend between dictation experiences and their opinion about it being positive or negative, perhaps due to our limited sample size.
5.4.2. On using Rambler for their own writing
Despite the learning curve of using speech for writing, 10 out of 12 participants preferred Rambler over the baseline and envisioned themselves using it in the future. The participant feedback suggests that Rambler is a versatile tool well-suited for a variety of tasks. First, we have text editing tasks: four participants mentioned situations where it would be useful to edit or break down text using Rambler. One participant said that semantic zoom would be useful for digesting content more quickly. Another participant said, ”What makes Rambler like better for certain use cases is being able to have like this higher level control of ideas. So I really enjoyed using those macro tools… I feel like this process of just like moving that you feel like you can actually like move an idea… in an abstract sense. I really enjoyed that part of it” (P4). Then, we have spoken-language tasks: Five participants mentioned using Rambler for transcribing conversations, interviews, or lectures, or composing speeches, presentations, or scripts. Five participants mentioned using Rambler for brainstorming, describing it as a way to ”capture the essence of things.” Three participants mentioned using Rambler for blogging or journaling (e.g., low-stake/shorter tasks).
6. Discussion
6.1. Summary of Findings
We found that writing with speech can be better supported by Rambler compared to the state-of-the-art practice, which is a STT text editor with the help of ChatGPT. Rambler is shown to be an effective tool for supporting text composition and revision via speech. Specifically, it showed advantages in helping users review spoken compositions, organize their thoughts, encouraging and supporting iteration while enhancing user control. Furthermore, participants used Rambler in highly creative and diverse ways, suggesting that Rambler is a versatile tool that can be adapted to different writing styles and workflows. The most frequently used features were keyword selection and semantic zoom. Macro revision features were mostly well received. Some LLM-assisted operations were less commonly used depending on user strategies, but participants who used them found them to be helpful. Finally, writing with speech was considered to be a good way to get started and capture loose thoughts. Despite the learning curve for getting used to writing with speech, almost all participants in the study preferred Rambler over the baseline and envisioned themselves using it in the future.
6.2. Effects of Gist-based Interface Features
The concept of gist-based interface was suggested for spoken text in previous theoretical work (Mehra et al., 2023) and explored in a few existing interfaces (Dang et al., 2022; Jiang et al., 2023) although not in this exact term. In this work, we further frame this concept in the context of dictation interfaces for long-form writing. It is operationalized with a few features, including: gist extraction through semantic zooming and keyword spotting; macro revision through Ramble respeaking, Ramble reorganization, LLM-assisted Ramble merging, splitting, and custom transformation.
Our data showed that gist extraction features were most frequently used. Keywords appeared to be highly versatile: visual grounding, users’ self-annotation, customizing summaries and influencing custom LLM transformation to align content with users’ voice. Semantic zoom had mixed ratings and was actively employed by fewer participants despite being tried by most. When employed, it was used in multiple ways including supporting review, serving as a checklist of ideas and generating concise versions of participants’ writing to be included in their final draft. Macro revision in essence refers to interaction with text in larger chunks, in order to support the conceptual level of text manipulation and revision. In theory, it should serve well for speech-to-text input, considering the speed of speech input, weaker verbatim memory, good contextual correction of ASR, and the difficulty of performing small corrections on mobile. In our study, manual operations for macro revisions were well-used, including respeaking and manual splitting / merging. While LLM-assisted splitting and merging were less used, participants came up with creative strategies to appropriate these features. Uncertainty about the outcome of these features remains an obstacle and affects users’ trust. The popularity and high rating of magic custom prompt reflects the need for supporting personalized requests in AI-assisted word processing.
Macro revision should go hand in hand with micro revision, as they complement each other in serving users’ goals.
Overall, Rambler sets a starting point for exploring the design space of gist-based interfaces for text. Our findings suggest that such interfaces can support useful GUI manipulations without always requiring high input precision.
6.3. Challenges Embedding LLM API Calls in a Latency-Sensitive GUI
Under the hood, many of Rambler’s system architecture decisions were driven by the latency capabilities and limitations of the underlying LLM (GPT-4) and transcription (AssemblyAI) APIs. We found GPT-4’s latency to be roughly linear to output length; for example, producing a 60-word summary took about 7 seconds.
Given that this would impact user experience in features like semantic zoom, we opted to pre-generate all summaries in parallel (and cache them for later) as soon as dictation ended, rather than waiting until users sought them out. We used the streaming form of the GPT-4 API to receive summaries. This approach showed users at least some summary text right away at each level, even if it would take several seconds to fully complete, improving interactivity.
As an illustration of how latency limitations and LLM capabilities interact to steer interfaces and interactions, consider the following anecdote from our development process: We initially attempted a mechanism for summary levels to update in real-time as the user was still dictating text, by sending that dictated-text-in-progress through our GPT-4 summary pipeline. However, this pipeline yielded a new summary with nearly zero word-level consistency with the old summary, even if the difference between new and old transcript was minor. A summary that completely changes every few seconds is not useful, but we found a way to specifically prompt GPT-4 to build on the previous existing summary based on the changes to the transcribed text from one summary requested to the next. Though this approach now generated useful “append-style” summaries, the prompts grew three times as fast as the transcript did—rapidly reaching latencies above 20 seconds, a delay so long as to render them essentially useless.
7. Limitations
One limitation of this work is that the study was conducted in a lab with an assigned task to participants. This artificial setting does not strongly motivate participants to feel invested in the task outcome. Given the nature of Rambler, this poses a challenge as participants are more likely to claim completion rather than continue editing, limiting the data we can collect. Although we designed the task specifically to try to mitigate this, there are likely still differences between what we could find in a 90 minute long session where participants were given a topic to write, in comparison with a more realistic task with greater benefit for the participant, in which the participant would feel more invested in the outcome. Additionally, a relatively small sample size leads to challenges of concluding some of the potential correlations, e.g. potential effects of participants’ dictation experiences.
Further, our choice of study device was an iPad tablet, representing one of the mobile platforms suitable for long-form writing. We believe our general concept of supporting writing with speech via rambles, gist extraction and macro revision can be applied to other mobile devices such as smartphones. However, a smaller screen would demand some UI adjustments. Not all our findings could directly generalize to a smartphone platform. For instance, the feature usages may differ as a smaller screen renders more challenges in content reviewing and manual editing, which points to a potentially greater need for support. Our current system already offers a responsive web interface, we plan to make UI adjustments and run future studies on other mobile devices in particular smartphones.
Rambler itself, as a research prototype, also comes with certain limitations. Firstly, Rambler semantic features relying on the GPT-4 API experience some unavoidable latency in the order of seconds. Our reliance on large language models for operations like semantic split and semantic merge might increase complications about proper attribution of authorship. Lastly, Rambler does not yet support multiple files or syncing across devices.
8. Future Work
Promising directions for future work include a longitudinal, diary study, which would provide participants with Rambler for longer and in a more realistic setting with their own mobile devices, potentially yielding insights into use less skewed by novelty effects. This diary study should also include a more diverse participant pool, addressing the limitation of our study being focused on an academic context. While Rambler showed a decent degree of versatility by surfacing highly diverse participant strategies in our lab study, future work shall investigate how to further support these diverse strategies, perhaps by making the features modular and user-customizable by choosing their own combinations. Future development of the approach described here could, for example, explore how to support personalization of LLM-powered GUIs through user-defined components for users’ own LLM operations. Lastly, here we merely scratched the surface of the design space of possible Gist-based Interfaces, supporting only a subset of the possible semantic operations made possible by LLMs. Extending this work to other use cases and other form factors beyond writing with speech on mobile devices could, in combination with this work, shed substantial light on new ways of writing more broadly.
Last but not least, we chose not to evaluate our participants’ text output via objective assessment and relied on self-reported measures for text quality. Our primary goal of the comparative study was to assess how well Rambler can provide users with control over the conceptual iteration of the text. As noted in §5.3.3, participants discussed retaining their writing style as a criterion while using ChatGPT or our LLM-assisted features. Future work could explore better ways to understand the nuances there. A standard external assessment of text quality, where we would recruit evaluators to rate it, would not work well in our study. Since LLMs are used in both Rambler and the baseline, the output almost always reads well and the quality measure is confounded with LLMs’ contribution. Existing methods for external assessment of text quality look primarily at attributes like language fluency or wording. In the age of GenAI, such methodologies shall be revisited and developed to account for such effects. Perhaps future research could include textual analysis examining text quality measures beyond cosmetic improvements, and focus more on conceptual development.
9. Conclusion
We designed, implemented, and evaluated Rambler, a versatile gist-based user interface that supports users to iteratively transform spontaneous rambly text into well-structured writing. With Rambler, users can start writing easily and quickly by rambling ideas into the interface, seeing the content in clean text, grasping its gists with semantic zooming and spotted keywords, and iterating on them via LLM-assisted macro revisions (i.e., respeaking, merging, splitting, and rearranging). Our evaluation showed major advantages of Rambler compared to an STT editor with ChatGPT, specifically in supporting reviewing, manipulating and revising the spoken text. The supported macro revision interactions are especially suitable for mobile devices as they relax the need for input precision. Our work also shows, in the context of writing with speech, injecting LLMs capability into a GUI tailored for its task can outperform a one-size-fit-all conversational interface of LLMs.
Appendix A Appendix
A.1. Timeline of Participant Workflows
Figure 10. Timelines visualizing participants’ workflow of each task using all the functionalities. The icons are manually annotated based on the coding of screen-recording videos. This reflects the qualitative observation of feature uses, consecutive triggers of functions may be annotated as single uses, and obvious accidental triggers for keyword highlighting were omitted.
A.2. List of Rambler Prompts
As mentioned in Section 3.4, we define custom prompts for our LLM-enabled functionality and embody them within our UI elements for ease of use. We list these prompts below.
A.2.1. Split Text Prompt
System: Split the paragraph the user enters into logical, cohesive paragraphs and return the result as a JSON array. Analyze the content and break it up into at least two separate paragraphs (but more where it makes sense). Try to split it into the appropriate number of paragraphs based on the content. Add each paragraph as a separate string element in the JSON array.
Response format: [”Paragraph 1 text”, ”Paragraph 2 text”, ”Paragraph 3 text”]
User: [User Ramble text]
A.2.2. Merge Text Prompt
System: You are a paragraph merger bot, capable of merging paragraphs. Please merge the following text into one paragraph of roughly median length as the originals:
You may use the following keywords to help you merge the text. Ensure that each keyword is in the merged paragraph.
The keywords are: [list of keywords].
Again, the resulting paragraph should be roughly the average length of the original paragraphs.
A.2.3. Clean Text Prompt
System: You are a text cleaning bot that cleans up the text the user enters by correcting obviously incorrect punctuation and formatting, but otherwise keeping the user’s text the exact same. You never ask your own questions. For example, if the user enters:
Google. Followed by the neural style transfer method, which became really popular. In a lot of cell phone apps. And these things started to. Show up in art galleries and art exhibitions. And more and more artists start playing with it. And now there’s a number of fairly. Um, significant contemporary artists who are also playing, experimented with AI techniques. This work all comes out of the, the academic research literature, and whatnot. So these are images from one academic paper which has really
you should return:
Google. Followed by the neural style transfer method, which became really popular in a lot of cell phone apps. And these things started to show up in art galleries and art exhibitions and more and more artists start playing with it. And now there’s a number of fairly, um, significant contemporary artists who are also playing, experimented with AI techniques. This work all comes out of the, the academic research literature, and whatnot. So these are images from one academic paper which has really
User: [User Ramble text]
A.2.4. Summarize Text Prompt
System: You are a professional writer specializing in text summarization. Make a summary of [LEVEL TEXT] of the chunk of the text provided by the user. The summary should reflect the main idea and the most important relationships of the text. You must preserve the same point of view, grammar and tense as the original text. If the text is in the first person, using words like I, you must use the first person as well. If the tone was conversational, you must be human conversational as well. You should use the following keywords to help you determine what to focus the summary on. Ensure that each keyword is in the summary. Try to fit as many as makes sense. Do not include anything else in the response other than the summary.
The keywords are: [list of keywords].
User: [User Ramble text]
where the possible values of [LEVEL TEXT], given that $L=$ the number of words in the user paragraph, are
  • 5 words or less
  • $L$ / 4 words or less
  • $L$ / 2 words or less